# Bias Amplification: Large Language Models as Increasingly Biased Media

### v2

@@ Appendix D Percentage of Center (Left) Biased Articles
Figure 7: This figure shows the percentage of center-leaning articles across multiple generations, comparing the baseline setup (’Synthetic’) with three mitigation strategies.
@@
Figure 8: This figure shows the percentage of left-leaning articles across multiple generations, comparing the baseline setup (’Synthetic’) with three mitigation strategies.
Appendix E Qualitative Bias Analysis Framework and Example of Bias Amplification Across Generations
@@
Story Framing and Selection Bias: This type of bias emerges when inherent leanings are found in the way topics, arguments, or narratives are structured. For instance, some aspects of reality are highlighted while others are obscured, shaping how the audience understands and interprets the events or issues at hand (Entman, 1993; Groeling, 2013). In extreme cases, opposing viewpoints are entirely excluded, leading to a one-sided representation of the issue. This selective omission restricts the audience’s comprehension of the full spectrum of perspectives, resulting in a distorted portrayal of the issue (Rodrigo-Ginés et al., 2024; Groeling, 2013). Entman described this as the selection and salience of specific facts that promote particular definitions, evaluations, and recommendations.
@@
Loaded Language Bias: This bias is identified through the use of charged or emotive words that signal political or ideological leanings. A common example is the difference in connotation between terms such as "undocumented" versus "illegal" immigrants. Such language choices often shape the audience’s perception by evoking specific emotional responses (Rodrigo-Ginés et al., 2024; Groeling, 2013).
@@
Below is an example of GPT-2 text outputs influenced by iterative synthetic training. The original article, titled "First Read: Why It’s So Hard for Trump to Retreat on Immigration, is a political opinion piece from NBC News, a left-leaning outlet as rated by AllSides (NBC News, 2016; AllSides, 2024b). The analysis follows the qualitative framework.
@@
  • Original Article: Why Its So Hard for Trump to Retreat on Immigration First Read is a morning briefing from Meet the Press and the NBC Political Unit on the day’s most important political stories and why they matter. Why its so hard for Trump to retreat on immigration Since launching his presidential candidacy 14 months ago, Donald Trumps most consistent and uncompromising policy issue has been immigration. Indeed, it was the subject of his first general-election TV ad that started airing on Friday. Yet over the weekend, his top aides and advisers suggested that Trump might be shifting on his past position that all of the 11 million undocumented immigrants living in the United States must be deported forcibly. To be determined, is what newly minted Campaign Manager Kellyanne Conway said on CNN when asked if Trump was retreating on the deportation force he talked about during the primary season. But here’s why its so hard – if not impossible – for Trump to retreat on immigration: Hes caught between his clear, unambiguous past statements and a base that might not willing to see him moderate on the issue. His past statements: Aug. 16, 2015 ""We’re going to keep the families together, but they have to go,"" Trump said on NBCs Meet the Press. More Trump: ""We will work with them. They have to go. Chuck, we either have a country, or we don’t have a country,"" he said. Nov. 11, 2015 You are going to have a deportation force, and you are going to do it humanely, Trump said on MSNBCs Morning Joe when asked how he would round up the nations 11 million undocumented immigrants. April 21, 2016 Look, were either going to have a country or were not going to have a country. But many people are very fine people. And I’m sure these are very, very fine people. They’re going to go, and were going to create a path where we can get them into this country legally, okay? But it has to be done legally – when asked by a questioner at a Today town hall that persons undocumented relatives would have to be deported if Trump becomes president. Trump cant ignore a base that has cheered his uncompromising immigration position And then there are the Trump supporters who’ve cheered the GOP presidential nominee for being so uncompromising on immigration. Classification Probability: 0.9946 for left-leaning, 0.0051 for center-leaning, 0.0002 for right-leaning
    Analysis of Story Framing and Selection Bias:
@@
    Omission of Alternative Perspectives. There is a lack of discussion about the potential benefits of Trump’s immigration policies or the concerns that might have led to his hardline approach. The article does not present viewpoints from supporters who agree with his stance or any arguments that justify his policies. This omission restricts the audience’s understanding of the full range of perspectives on the issue
@@
  • Generation 0: Why Its So Hard for Trump to Retreat on Immigration The president’s decision to withdraw from the Paris climate accord is a major blow to the United States’ standing in the world. The United States has been a major player in the world of climate change, and has been a major player in the fight against climate change. But Trump’s decision to immigration. He has repeatedly said that he would deport all illegal immigrants, but has repeatedly said that he would not deport all illegal immigrants. He has also said that he would not deport all illegal immigrants who have committed crimes, but has repeatedly said that he would not deport all illegal immigrants who have committed crimes. He has also clear, this is not a new position. It was a position that Trump has held for years. But it is a position that he has held for years. And it is a position that he has held for years. And it is a position that he has held for years. And it is a position that he has the deportation force. 1. Trump is not retreating on the deportation force. Trump is not retreating on the deportation force. Trump is not retreating on the deportation force. Trump is not retreating on the the Press. "We’re going to keep the families together, but they have to go. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together.ly. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely We’re going to have a country that is going to be a great place for the American people. We’re going to have a country that is going to be a great place for the American people. We’re going to have a country that is going to be a great place for the American people. Classification Probability: 0.0011 for left-leaning, 0.9957 for center-leaning, 0.0032 for right-leaning
    Analysis of Story Framing Bias and Selection:
@@
  • Generation 4: Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on Immigration, and why its so hard for Trump to Retreat on immigration, and the presidents speech was a stark reminder that the president is not the presidents speech was a stark reminder that the president is not the economy the presidents campaign manager, Kellyanne Conway, said the presidents comments were a complete fabrication. The presidents comments are a complete fabrication. Conway said the presidents comments were a complete fabrication. a, the presidents speech was a stark reminder that the president is not a man of words. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a the issues campaign manager, Kellyanne Conway, said the campaign is actively working with the FBI to determine the identity of the person who leaked the emails. We are working with the FBI to determine the identity of the person who leaked the emails, Conway said in a statement the Press the presidents comments. I think its a very, very sad day for the country, Trump said on Fox News Sunday. I think its a very, very sad day for the country for the country for the country forly. The presidents speech was a stark reminder that the president is not a man of words. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a the presidents speech was a stark reminder that the president is not a politician. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a man of to the the presidents executive actions on immigration. The presidents order, which was signed into law by President Barack Obama on Friday, suspends the entry of refugees and travelers from seven majority-Muslim countries, including Iran, Iraq, Libya, Somalia, Sudan, Syria and Yemen. Classification Probability: 0.0006 for left-leaning, 0.0044 for center-leaning, 0.9950 for right-leaning
    Analysis of Story Framing and Selection Bias:
@@
  • Generation 10: Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on immigration. The presidents speech was a stark reminder that the president the president the president the presidents the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the the the president the president the president the president the president the president the presidents statement that the president has not yet made a decision on whether to fire Comey. The president has not yet made a decision on whether to fire Comey, Mr. Trump the Press the president the president the president the president the president the president the presidents statement that the president has not yet made a decision on whether to fire Comey. The president has not yet made a decision on whether to fire Comey, Mr. Trumply the the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he The the the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he said the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were Classification Probability: 0.0073 for left-leaning, 0.4127 for center-leaning, 0.5800 for right-leaning
    Analysis of Story Framing and Selection Bias:
@@
Appendix F Distribution of Text Quality Index
Figure 9: This figure presents the distribution of the text quality index across multiple generations for the baseline setup.
Appendix G Average Perplexity Across Generations
Figure 10: This figure shows the average perplexity for synthetic articles , along with 95% confidence intervals, across multiple generations for the baseline setup.
Appendix H Example of Quality Deterioration Across Generations
Examples of GPT-2 text outputs affected by iterative synthetic (Example articles are truncated for brevity, as the full articles are lengthy but exhibit consistent quality degradation throughout; see the dataset for details).
H.1 Example of Deterministic Setup
  • Original Article: The world’s eyes are on President Obama this week as he rallies a coalition of nations to “destroy” the extremist Islamic State (IS) and its terrorist-led territory. Yet if the Arab world is ever to throw off its many forms of tyranny – from theocracy to autocracy to monarchy – it also needs a model to emulate. Text Quality Index: 2.81
  • Generation 0: , the Syrian government has launched a series of airstrikes on the militant group’s stronghold of Raqqa, the capital of the self-proclaimed Islamic State. The strikes have targeted the Islamic State’s military headquarters, the Al-Bab, a military training center, and the Al-Baba, a military training center in it will need to confront the Islamic State’s growing influence in the region. Text Quality Index: 2.58; Perplexity: 6.68
  • Generation 4: in Iraq and Syria (ISIS) group Read more The Iraqi army has been fighting the Islamic State since the group seized large swaths of territory in Iraq and Syria (ISIS) group in 2014. The Iraqi army has been fighting the Islamic State the Iraqi army. The move comes as the U.S. the Iraqi the the the the holiest places in the world. Text Quality Index: 2.01; Perplexity: 3.17
  • Generation 10: the Iraqi the Iraqi army. The move comes as the United States and its allies are ramping up their military campaign against the Islamic State, the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi army. The the Iraqi the the the the holiest the holiest the holiest the holiest the holiest places in the world. The attack came just hours after a suicide bomber blew himself up at a Christmas market in Nice, killing at least 32 people and injuring scores more. Text Quality Index: 1.24; Perplexity: 4.23
Appendix I Mathematical Details for the Statistical Tests
We will explain how the relationship between changes in neuron weight and changes in bias performance (generation quality) can be statistically tested.
@@ Appendix I Mathematical Details for the Statistical Tests
Appendix J Literature Review of Model Collapse
Model Collapse. Shumailov et al. (2024); Alemohammad et al. (2023); Guo et al. (2024); Wyllie et al. (2024); Dohmatob et al. (2024a) describe it as a degenerative process in which models, recursively fed with their own data, increasingly distort reality and lose generalizability, for example, by prioritizing high-probability events while neglecting rare ones or shifting distributions. Shumailov et al. (2024), utilizing the OPT-125M, demonstrates this phenomenon, showing that the perplexity distribution becomes increasingly skewed, with more concentration at lower perplexities and longer tails. Taori and Hashimoto (2022) observed an increase in repetitive content during the synthetic fine-tuning of GPT-2. Similarly, Guo et al. (2024); Dohmatob et al. (2024b); Seddik et al. (2024) show that the OPT-350M, Llama2, and GPT2-type models experience performance deterioration after several generations, such as a decrease in linguistic diversity or greater divergences in token probabilities. Alemohammad et al. (2023) studies model collapse in generative image models, finding that quality and diversity deteriorate with synthetic training loops. They also found that cherry-picking high-quality outputs by users contributes to sampling errors, which actually helps maintain quality. Interestingly, Hamilton (2024) found that GPT-3.5-turbo shows less diversity in perspectives in narrative writing tasks compared to earlier models like davinci-instruct-beta and text-davinci-003.

### v3

@@ Appendix B Percentage of Center (Left) Biased Articles
Figure 6: Evolution of center-leaning article percentage across generations, comparing baseline (’Synthetic’) with three mitigation strategies (Main Experiment).
@@
Figure 7: Evolution of left-leaning article percentage across generations, comparing baseline (’Synthetic’) with three mitigation strategies (Main Experiment).
Appendix C Qualitative Bias Analysis
@@
Story Framing and Selection Bias. This type of bias emerges when inherent leanings are found in the way topics, arguments, or narratives are structured. For instance, some aspects of reality are highlighted while others are obscured, shaping how the audience understands and interprets the events or issues at hand (Entman, 1993; Groeling, 2013). In extreme cases, opposing viewpoints are entirely excluded, leading to a one-sided representation of the issue. This selective omission restricts the audience’s comprehension of the full spectrum of perspectives, resulting in a distorted portrayal of the issue (Rodrigo-Ginés et al., 2024; Groeling, 2013). Entman described this as the selection and salience of specific facts that promote particular definitions, evaluations, and recommendations.
@@
Loaded Language Bias. This bias is identified through the use of charged or emotive words that signal political or ideological leanings. A common example is the difference in connotation between terms such as "undocumented" versus "illegal" immigrants. Such language choices often shape the audience’s perception by evoking specific emotional responses (Rodrigo-Ginés et al., 2024; Groeling, 2013).
@@
Below is an example of GPT-2 text outputs influenced by iterative synthetic training. The original article, titled "First Read: Why It’s So Hard for Trump to Retreat on Immigration, is a political opinion piece from NBC News, a left-leaning outlet as rated by AllSides (NBC News, 2016; AllSides, 2024b). The analysis follows the qualitative framework:
@@
Original Article. Why Its So Hard for Trump to Retreat on Immigration First Read is a morning briefing from Meet the Press and the NBC Political Unit on the day’s most important political stories and why they matter. Why its so hard for Trump to retreat on immigration Since launching his presidential candidacy 14 months ago, Donald Trumps most consistent and uncompromising policy issue has been immigration. Indeed, it was the subject of his first general-election TV ad that started airing on Friday. Yet over the weekend, his top aides and advisers suggested that Trump might be shifting on his past position that all of the 11 million undocumented immigrants living in the United States must be deported forcibly. To be determined, is what newly minted Campaign Manager Kellyanne Conway said on CNN when asked if Trump was retreating on the deportation force he talked about during the primary season. But here’s why its so hard – if not impossible – for Trump to retreat on immigration: Hes caught between his clear, unambiguous past statements and a base that might not willing to see him moderate on the issue. His past statements: Aug. 16, 2015 ""We’re going to keep the families together, but they have to go,"" Trump said on NBCs Meet the Press. More Trump: ""We will work with them. They have to go. Chuck, we either have a country, or we don’t have a country,"" he said. Nov. 11, 2015 You are going to have a deportation force, and you are going to do it humanely, Trump said on MSNBCs Morning Joe when asked how he would round up the nations 11 million undocumented immigrants. April 21, 2016 Look, were either going to have a country or were not going to have a country. But many people are very fine people. And I’m sure these are very, very fine people. They’re going to go, and were going to create a path where we can get them into this country legally, okay? But it has to be done legally – when asked by a questioner at a Today town hall that persons undocumented relatives would have to be deported if Trump becomes president. Trump cant ignore a base that has cheered his uncompromising immigration position And then there are the Trump supporters who’ve cheered the GOP presidential nominee for being so uncompromising on immigration.
Classification Probability: 0.9946 for left-leaning, 0.0051 for center-leaning, 0.0002 for right-leaning.
Analysis:
@@
Omission of Alternative Perspectives. There is a lack of discussion about the potential benefits of Trump’s immigration policies or the concerns that might have led to his hardline approach. The article does not present viewpoints from supporters who agree with his stance or any arguments that justify his policies. This omission restricts the audience’s understanding of the full range of perspectives on the issue.
@@
Generation 0. Why Its So Hard for Trump to Retreat on Immigration The president’s decision to withdraw from the Paris climate accord is a major blow to the United States’ standing in the world. The United States has been a major player in the world of climate change, and has been a major player in the fight against climate change. But Trump’s decision to immigration. He has repeatedly said that he would deport all illegal immigrants, but has repeatedly said that he would not deport all illegal immigrants. He has also said that he would not deport all illegal immigrants who have committed crimes, but has repeatedly said that he would not deport all illegal immigrants who have committed crimes. He has also clear, this is not a new position. It was a position that Trump has held for years. But it is a position that he has held for years. And it is a position that he has held for years. And it is a position that he has held for years. And it is a position that he has the deportation force. 1. Trump is not retreating on the deportation force. Trump is not retreating on the deportation force. Trump is not retreating on the deportation force. Trump is not retreating on the the Press. "We’re going to keep the families together, but they have to go. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together.ly. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely We’re going to have a country that is going to be a great place for the American people. We’re going to have a country that is going to be a great place for the American people. We’re going to have a country that is going to be a great place for the American people.
Classification Probability: 0.0011 for left-leaning, 0.9957 for center-leaning, 0.0032 for right-leaning.
Analysis:
@@
Generation 4. Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on Immigration, and why its so hard for Trump to Retreat on immigration, and the presidents speech was a stark reminder that the president is not the presidents speech was a stark reminder that the president is not the economy the presidents campaign manager, Kellyanne Conway, said the presidents comments were a complete fabrication. The presidents comments are a complete fabrication. Conway said the presidents comments were a complete fabrication. a, the presidents speech was a stark reminder that the president is not a man of words. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a the issues campaign manager, Kellyanne Conway, said the campaign is actively working with the FBI to determine the identity of the person who leaked the emails. We are working with the FBI to determine the identity of the person who leaked the emails, Conway said in a statement the Press the presidents comments. I think its a very, very sad day for the country, Trump said on Fox News Sunday. I think its a very, very sad day for the country for the country for the country forly. The presidents speech was a stark reminder that the president is not a man of words. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a the presidents speech was a stark reminder that the president is not a politician. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a man of to the the presidents executive actions on immigration. The presidents order, which was signed into law by President Barack Obama on Friday, suspends the entry of refugees and travelers from seven majority-Muslim countries, including Iran, Iraq, Libya, Somalia, Sudan, Syria and Yemen.
Classification Probability: 0.0006 for left-leaning, 0.0044 for center-leaning, 0.9950 for right-leaning.
Analysis:
@@
Generation 10. Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on immigration. The presidents speech was a stark reminder that the president the president the president the presidents the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the the the president the president the president the president the president the president the presidents statement that the president has not yet made a decision on whether to fire Comey. The president has not yet made a decision on whether to fire Comey, Mr. Trump the Press the president the president the president the president the president the president the presidents statement that the president has not yet made a decision on whether to fire Comey. The president has not yet made a decision on whether to fire Comey, Mr. Trumply the the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he The the the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he said the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were
Classification Probability: 0.0073 for left-leaning, 0.4127 for center-leaning, 0.5800 for right-leaning.
Analysis:
@@
Appendix D Distribution of Text Quality Index
Figure 8: Distribution of Text Quality Index across generations for the baseline experiment (no mitigation), showing progressive degradation.
Appendix E Average Perplexity Across Generations
Figure 9: Evolution of average perplexity (95% CIs) across generations for the baseline experiment (no mitigation).
Appendix F Example of Quality Deterioration Across Generations
Examples of GPT-2 text outputs affected by iterative synthetic (Example articles are truncated for brevity).
Original Article. The world’s eyes are on President Obama this week as he rallies a coalition of nations to "destroy" the extremist Islamic State (IS) and its terrorist-led territory. Yet if the Arab world is ever to throw off its many forms of tyranny – from theocracy to autocracy to monarchy – it also needs a model to emulate.
Text Quality Index: 2.81
Generation 0. the Syrian government has launched a series of airstrikes on the militant group’s stronghold of Raqqa, the capital of the self-proclaimed Islamic State. The strikes have targeted the Islamic State’s military headquarters, the Al-Bab, a military training center, and the Al-Baba, a military training center in it will need to confront the Islamic State’s growing influence in the region.
Text Quality Index: 2.58; Perplexity: 6.68
Generation 4. in Iraq and Syria (ISIS) group Read more The Iraqi army has been fighting the Islamic State since the group seized large swaths of territory in Iraq and Syria (ISIS) group in 2014. The Iraqi army has been fighting the Islamic State the Iraqi army. The move comes as the U.S. the Iraqi the the the the holiest places in the world.
Text Quality Index: 2.01; Perplexity: 3.17
Generation 10. the Iraqi the Iraqi army. The move comes as the United States and its allies are ramping up their military campaign against the Islamic State, the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi army. The the Iraqi the the the the holiest the holiest the holiest the holiest the holiest places in the world. The attack came just hours after a suicide bomber blew himself up at a Christmas market in Nice, killing at least 32 people and injuring scores more.
Text Quality Index: 1.24; Perplexity: 4.23
Appendix G Mathematical Details for the Statistical Tests
We now explain how the relationship between changes in neuron weights and changes in bias performance (or generation quality) can be statistically tested.
First, we compute the test statistic as $t_{\beta_{j}}=\frac{\beta_{j}}{SE(\beta_{j})}$, where $SE(\beta_{j})$ is the standard error of $\beta_{j}$, estimated using the Newey–West estimator to account for potential heteroscedasticity and autocorrelation in the residuals.
@@ Appendix G Mathematical Details for the Statistical Tests
Appendix H Literature Review of Mitigation Strategies